HUBBLE SPACE TELESCOPE DAILY REPORT #1050 PERIOD COVERED: 0700 EST Daily Status Report as of 011/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5190 (Observations of External Galaxies) The WF/PC-2 was used to make images of a distant cluster of galaxies with a background quasi-stellar object. A total of seven images were received and displayed. There were many recentering events reported during each of the first five exposures. The last two images were reported to be okay. 1.2 Completed WF/PC-2 4709 (WF/PC-2 Internal Monitor) The WF/PC-2 was used to make a set of routine internal monitoring exposures. These include K-spots and and internal flat images. A total of 20 images were received and displayed. They were reported to look okay. 1.3 Completed WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibration) The WF/PC-2 was used to make a set of visual internal flat images. Six images were received and displayed. They were reported to be okay. One image still remains on the science tape recorder, and will be analyzed later. 1.4 Completed FOC/96 4932 (Post-COSTAR FOC/96 Fine Aperture Location) The Faint Object Camera was used to make images of cluster NGC-188 to measure the location of the post-COSTAR f/96 aperture with respect to the V2-V3 coordinate frame. Two images were taken, and no anomalies were noted. 1.5 Completed FGS 4892 (Duplicity Among Stars in an External Cluster, Cycle 3 High) The FGSs were used to make astrometric observations of the target star UPGREEN-69 located in the Hyades open cluster. The FGS were monitored during the scans, and no problems were seen in the telemetry. 1.6 Completed FOC/96 5191 (High Resolution Ultraviolet Observations of an External Source) The Faint Object Camera (f/96) was used to make deep ultraviolet images of the globular cluster M-15 (aka, NGC-7078). The first image was received in real time and displayed in the OSS. A corrective SIP maneuver was successfully uplinked to the spacecraft. Subsequent observations were reported to be as expected. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled :8 Successful :8 2.3 Significant HSTARs: None 2.4 Operations Notes: All SMOV activities are being performed as planned, and the status of all onboard subsystems continue to be nominal. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibration) 3.2 Complete FOS/BL 4907 (SMOV FOS/FGS Coarse Alignment) 3.3 Complete WF/PC-2 4616 (WF/PC-2 Fine Alignment and Focus) 3.4 Complete FOC/96 4787 (SMOV FOC Throughput and Sensitivity) 3.5 Complete WF/PC-2 5190 (Observations of External Galaxies) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-